OpenStudy (anonymous):

Factor 5t^2+14t+8 or is it prime ?

OpenStudy (anonymous):

you could always try the quadratic formula to check

OpenStudy (anonymous):

Whats that ? How ?

OpenStudy (anonymous):

but it appears to be prime

OpenStudy (anonymous):

-b+or- sq.rt. of (b^2-4ac) all divided by 2a

OpenStudy (anonymous):

a=coefficient of t^2 b=coefficient of t c=the constant

OpenStudy (anonymous):

\[5 t^2+14 t+8=(t+2) (5 t+4) \]

OpenStudy (anonymous):

no its not robto is right

OpenStudy (anonymous):

- 0HKAY THANKS .